Riverbed Company is a multi product firm. Presented below is information concerning one of its products, the Hawkeye.

Date

Transaction

Quantity

Price/Cost

1/1 Beginning inventory 3,200 $18
2/4 Purchase 4,200 27
2/20 Sale 4,700 46
4/2 Purchase 5,200 35
11/4 Sale 4,400 50

1.

Compute cost of goods sold, assuming Riverbed uses: (Round average cost per unit to 4 decimal places, e.g. 2.7631 and final answers to 0 decimal places, e.g. 6,548.)

Cost of goods sold
(a) Periodic system, FIFO cost flow $

(b) Perpetual system, FIFO cost flow $

(c) Periodic system, LIFO cost flow $

(d) Perpetual system, LIFO cost flow $

(e) Periodic system, weighted-average cost flow $

(f) Perpetual system, moving-average cost flow $

Please show all work.
Answer #1

ANSWER

Cost of goods available for sale = (3200 * $18) + (4200*$27) + (5200*$35) = $353,000

Total quantity available for sale = 3200 + 4200 +5200 = 12600

Average cost per unit = Cost of goods available for sale / Nos of units available for sale

= $353,000 / 12600 = $28.02 per unit

Solution a:

Cost of goods available for sale = (3200 * $18) + (4200*$27) + (5200*$35) = $353,000

Quantity in ending inventory = 3200 + 4200 +5200 - 4700 - 4400 = 3500

In FIFO ending inventory will consist of purchase from 4/2

Hence value of ending inventory = 3500 * $35 = $122,500

Cost of goods sold = Cost of goods available for sale - Value of ending inventory

= $353,000 - $122,500 = $230,500

Solution b:

Computation of ending inventory COGS under FIFO
Date Beginning Inventory Purchase Cost of Goods Sold Ending Inventory
Qty Rate Amount Qty Rate Amount Qty Rate Amount Qty Rate Amount
1/1 3200 $18.00 $57,600.00 0 $0.00 $0.00 0 $0.00 $0.00 3200 $18.00 $57,600.00
2/4 3200 $18.00 $57,600.00 4200 $27.00 $113,400.00 0 $0.00 $0.00 3200 $18.00 $57,600.00
4200 $27.00 $113,400.00
2/20 3200 $18.00 $57,600.00 0 $0.00 $0.00 3200 $18.00 $57,600.00 2700 $27.00 $72,900.00
4200 $27.00 $113,400.00 1500 $27.00 $40,500.00
4/2 2700 $27.00 $72,900.00 5200 $35.00 $182,000.00 0 $0.00 $0.00 2700 $27.00 $72,900.00
5200 $35.00 $182,000.00
11/4 2700 $27.00 $72,900.00 0 $0.00 $0.00 2700 $27.00 $72,900.00 3500 $35.00 $122,500.00
5200 $35.00 $182,000.00 1700 $35.00 $59,500.00
Total $230,500.00 $122,500.00

Solution c:

Under LIFO ending inventory will consist of 3200 units from begining inventory and 300 units from 2/4 purchase

Cost of ending inventory = (3200*$18) + (300*$27) = $65,700

Cost of goods sold = $353,000 - $65,700 = $287,300

Solution d:

Computation of ending inventory COGS under LIFO
Date Beginning Inventory Purchase Cost of Goods Sold Ending Inventory
Qty Rate Amount Qty Rate Amount Qty Rate Amount Qty Rate Amount
1/1 3200 $18.00 $57,600.00 0 $0.00 $0.00 0 $0.00 $0.00 3200 $18.00 $57,600.00
2/4 3200 $18.00 $57,600.00 4200 $27.00 $113,400.00 0 $0.00 $0.00 3200 $18.00 $57,600.00
4200 $27.00 $113,400.00
2/20 3200 $18.00 $57,600.00 0 $0.00 $0.00 4200 $27.00 $113,400.00 2700 $18.00 $48,600.00
4200 $27.00 $113,400.00 500 $18.00 $9,000.00
4/2 2700 $18.00 $48,600.00 5200 $35.00 $182,000.00 0 $0.00 $0.00 2700 $18.00 $48,600.00
5200 $35.00 $182,000.00
11/4 2700 $18.00 $48,600.00 0 $0.00 $0.00 4400 $35.00 $154,000.00 2700 $18.00 $48,600.00
5200 $35.00 $182,000.00 800 $35.00 $28,000.00
Total $276,400.00 $76,600.00

Solution e:

Weighted average cost per unit = Cost of goods available for sale / Nos of units avaiable for sale

= $353,000 / 12600 = $28.02 per unit

Cost of ending inventory using Weighted average = 3500 * $28.02 = $98,056

Cost of goods sold =$353,000 - $98,056 = $254,944

Solution f:

Computation of ending inventory COGS under Weighted Average Cost
Date Opening Stock Receiving Issue Closing Stock
Qty Rate Amount Qty Rate Amount Qty Rate Amount Qty Rate Amount
1/1 3200 $18.00 $57,600.00 0 $0.00 $0.00 0 $0.00 $0.00 3200 $18.00 $57,600.00
2/4 3200 $18.00 $57,600.00 4200 $27.00 $113,400.00 0 $0.00 $0 7400 $23.11 $171,000
2/20 7400 $23.11 $171,000 0 $0.00 $0 4700 $23.11 $108,608 2700 $23.11 $62,392
4/2 2700 $23.11 $62,392 5200 $35.00 $182,000 0 $0.00 $0 7900 $30.94 $244,392
11/4 7900 $30.94 $244,392 0 $0.00 $0 4400 $30.94 $136,117 3500 $30.94 $108,275
Total $244,725 $108,275

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************